The budgeted income statement presented below is for Burkett Corporation for the coming fiscal year. If Burkett Corporation is able to achieve the budgeted level of sales, its margin of safety in dollars would be (Do not round intermediate calculations.):

Sales (58,000 units) $986,000

Costs: Direct materials $149,200

Direct labor 240,800

Fixed factory overhead 104,000

Variable factory overhead 150,800

Fixed marketing costs 110,800

Variable marketing costs 50,800 806,400

Pretax income $179,600.
Answer #1

Answer = $449,000

Margin of Safety = Actual dollar Sales - BEP Dollar Sales

Actual Dollar Sales = 986,000

BEP Dollar Sales = Fixed Cost / CM ratio

CM Ratio = Contribution Margin / Sales Revenue

Contribution Margin = Sales - Variable cost = 986,000-149200-240800-150800-50,800 = 394,400

CM Ratio = 394,400/986,000 = 40%

Break Even Dollars = (104,000+110,800)/40% =214,800 /40%= 537,000

Margin of Safety = 986,000-537,000 = $ 449,000 (Answer)